The pressure inside a tank of neon is 110 atm . The temperature is 28.0 ∘C.

On average, how many atomic diameters does a neon atom move between collisions?
Answer #1

PV = NKT

N/V = P/KT

N/V = 110*101325/1.38*10^-23*310 = 26053.6*10^23 m^-3

mean free path of neaon atom is 12.7*10^-3

mean free path = KT/1.414*3.14*P*D^2

D^2 = 1.38*10^-23*301/1.414*3.14*110*101325*12.7*10^-3

D = 8.13*10^-14 M

R = D/2 = 4.1*10^-14 m